### v2

# Zephyrus: Scaling Gateways Beyond the Petabit-Era with DPU-Augmented Hierarchical Co-Offloading

Abstract.
Operating at petabit-scale, ByteDance’s cloud gateways are deployed at critical aggregation points to orchestrate a wide array of business traffic. However, this massive scale imposes significant resource pressure on our previous-generation cloud gateways, rendering them unsustainable in the face of ever-growing cloud-network traffic. As the DPU market rapidly expands, we see a promising path to meet our escalating business traffic demands by integrating DPUs with our established Tofino-based gateways. DPUs augment these gateways with substantially larger table capacities and richer programmability without compromising previously low-latency and high-throughput forwarding. Despite compelling advantages, the practical integration of DPUs into cloud gateways remains unexplored, primarily due to underlying challenges.
In this paper, we present Zephyrus, a production-scale gateway built upon a unified P4 pipeline spanning high-performance Tofino and feature-rich DPUs, which successfully overcomes these challenges. We further introduce a hierarchical co-offloading architecture (HLCO) to orchestrate traffic flow within this heterogeneous gateway, achieving $99\%$ hardware offloading while retaining software fallback paths for complex operations. Zephyrus outperforms LuoShen (NSDI ’24) with 33% higher throughput and our evaluation further indicates 21% lower power consumption and 14% lower hardware cost. Against FPGA-based systems, Albatross (SIGCOMM ’25), it doubles the throughput at a substantially lower Total Cost of Ownership (TCO), showcasing its superior performance-per-dollar. Beyond these performance gains, we also share key lessons from several years of developing and operating Zephyrus at production scale. We believe these insights provide valuable references for researchers and practitioners designing performant cloud gateways.
conference: ; ; footnotetext: The first three authors contribute equally. Tong Yang (yangtong@pku.edu.cn) and Xiyun Xu (wumin.stephen@bytedance.com) are corresponding authors.
1. Introduction
As shown in Figure 1, ByteDance’s cloud and edge services generate petabit-scale aggregate traffic across tens of thousands of tenants. Orchestrating these workloads critically depends on cloud gateways, which sit at aggregation points and mediate intra-datacenter (east-west) and external (north-south) traffic (google:jupiter; google:maglev), including IDC and cross-region connectivity. As the linchpin of our network, gateways must simultaneously deliver low latency, sustain massive throughput, and remain cost efficient at scale (lu2025albatross; pan2024luoshen). These requirements make gateway one of the most performance-critical components in our infrastructure.
Diverse workloads drive the demand for high-performance gateways. On our cross-region LLM platform Doubao (doubao), user requests are first processed at nearby edge PoPs and then traverse backbone links to regional clusters, generating sustained cross-domain traffic through border gateways. On the short-video platform TikTok (tiktok), video clips are relayed across edge servers while real-time AI features (e.g., filters, face tracking, speech recognition) run at the edge and heavier tasks are handled in the cloud; this hybrid execution creates continuous, high-bandwidth flows between edge and cloud. On Volcengine (volcengine), our enterprise cloud platform, multi-region deployments synchronize storage and services across sites, driving frequent large-scale cross-region transfers, especially in hybrid on-premises–cloud settings. Together, these workloads place gateways on the critical path, requiring them to sustain massive throughput, low-latency forwarding, and scalable multi-tenant policies.
Figure 1. Aggregate traffic from ByteDance cloud and edge services
Our experience with two generations of cloud gateways, together with observations from prior deployments and research, highlights the fundamental limitations of existing approaches. Our first generation of CPU-based software gateways provides flexibility, but fails to scale with exponential traffic growth, suffering from per-core bottlenecks, performance isolation issues, and vulnerability to flow imbalance (lei2024enhancing; pismenny2021alleviating; kfoury2024smartnic). The second version of Tofino-based gateways delivers line-rate forwarding with microsecond latency, but faces hard resource ceilings. Although algorithmic LPM (ALPM) and pipeline folding (pan2024luoshen; pan2021sailfish) can be applied to expand capacity, production workloads still induced severe localized hotspots. Other works also explore the possibility of FPGA-based gateways, which offer greater flexibility and can implement relatively complex operations that exceed the limits of ASIC resources (lu2025albatross). However, FPGAs incur substantially higher hardware cost and power consumption compared to programmable ASICs (wang2020multitenancy; yang2024p4eip). Moreover, FPGA programming still relies on low-level HDLs (e.g., Verilog, VHDL) and complex vendor toolchains, resulting in much longer development cycles and higher overheads (sozzo2022survey; jamal2018fpga; windh2015fpga). Taken collectively, these limitations underscore the fundamental tension between performance, programmability, development difficulty, and cost efficiency in gateway design.
To overcome these limitations, we present our third generation of cloud gateways, Zephyrus. To the best of our knowledge, Zephyrus is the world’s first heterogeneous cloud gateway that integrates programmable switching ASICs with DPUs under a unified control plane. DPUs are the key enablers of Zephyrus: they work in concert with switching ASICs, achieving flexibility, easy development, and cost efficiency, while maintaining low-latency forwarding and high throughput. By relocating the VM location table to DPUs with abundant memory, Zephyrus alleviates the severe on-chip resource pressure of ASIC pipelines.
However, designing and deploying Zephyrus is a non-trivial task. First, Zephyrus requires precise coordination across heterogeneous components, where task partitioning, functional integration, effective communication between ASICs, DPUs, and software must be carefully orchestrated. Second, DPUs are relatively unexplored in the context of cloud gateways, which compelled us to address their integration complexity largely without existing design blueprints  (lu2025albatross). Last but not least, system upgrades at petabit scale require careful planning to ensure business continuity, smooth expansion, and robust disaster recovery.
To this end, we design a novel Hierarchical Co-Offloading (HLCO) architecture, which extends the classic fast/slow path model into three tiers: Switching ASICs handle high-throughput, low-complexity processing; DPUs provide a middle tier with greater flexibility; and software offers full programmability for most flexible processing, allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. For compatibility, we implemented Zephyrus on P4-programmable DPUs (Pensando (pensando2022)) and switching ASICs (Tofino), as both support P4 and thus enable a unified programming model. Four Pensando DPUs are integrated with a folded Tofino pipeline in our deployment, achieving 1.6 Tbps line-rate forwarding. To support efficient coordination, we carefully design a minimal yet expressive metadata format for communication between Tofinos and Pensando DPUs. In the control plane, we develop a version-based flow update mechanism to ensure synchronization of the forwarding rules. Beyond introducing a new gateway design, this paper shares the experiences and lessons distilled from several years of development and production deployment, providing insights that may inform both researchers and practitioners building next-generation cloud gateways.
Our major contributions are summarized as follows.
[…]
  • We propose a novel hierarchically co-offloading architecture (HLCO) that enables efficient partitioning of P4 data-plane processing across high-performance switching ASICs and DPUs. HLCO combines the strength of the classic fast/slow path model and flexible 3-stage offloading (li2024triton), allowing Zephyrus to harness the strengths of heterogeneous hardware while mitigating their weaknesses. With HLCO, more than 99% of packets are offloaded to hardware, minimizing software overhead while retaining the flexibility.
  • We evaluate Zephyrus against LuoShen (pan2024luoshen), a state-of-the-art programmable gateway. Zephyrus delivers 33% higher throughput and is estimated to reduce power consumption by at least 21% and hardware procurement cost by 14% per 2U chassis, under comparable configurations. Compared to Albatross (lu2025albatross), another state-of-the-art FPGA-based cloud gateway, Zephyrus achieves 2× higher throughput while requiring significantly lower estimated total cost of ownership (TCO), demonstrating a superior performance-per-dollar ratio.
2. Background and Motivation
Figure 2. Cloud network gateway architecture.
This section provides a context for our work. We first provide a brief overview and analysis of our cloud gateway, followed by its architectural evolution. We further analyze these challenges, which reveal the limitations of prior generations and ultimately motivate the architectural evolution towards Zephyrus.
2.1. Anatomy of Cloud Gateways
Roles of cloud gateways. As shown in Figure 2, the Cloud Gateway Architecture across Regions, cloud gateways are packet forwarding devices at the network boundary, serving as the linchpin of the cloud network by orchestrating all data flows (chiang2018andromeda). As the central hub, they manage major traffic patterns from east to west (VM-VM) to north to south (VM-cross-region-VM and VM-IDC) (greenberg2009vl2; dean2013tail) with their corresponding datapaths detailed in Table 1. In ByteDance, cloud gateways serve as the high-performance backbone for a wide range of upper-layer applications (e.g., distributed AI services, large-scale content delivery, and hybrid cloud interconnection), providing the low-latency, high-throughput forwarding required for their business-critical traffic.
Table 1. Major cloud services and corresponding traffic routes (CGW for could gateway)
Cloud services Traffic routes
VM–VM (the same region) VM–vSwitch–CGW–vSwitch–VM
VM–Cross-region–VM VM–vSwitch–CGW–Cross-region–CGW–vSwitch–VM
VM–IDC VM–vSwitch–CGW–Customer switch–IDC
Characteristics of gateway workloads. Gateway workloads exhibit several important characteristics. (1) A significant portion of traffic is highly latency-sensitive (barker2010empirical; wu2023optimizing). As mentioned above, critical applications such as LLM inference, real-time collaboration, and recommendation services impose strict latency limits, where even minor delays can degrade service quality and ultimately compromise service level objectives (SLO). (2) The traffic is characterized by its immense aggregate volume, with aggregate traffic reaching the petabit-per-second (Pbps) scale. This is a direct consequence of simultaneously serving numerous data-intensive applications, from large-scale content delivery to distributed AI services. (3) The traffic stream is inherently a composite of millions of concurrent flows (chen2014concurrent; bates2014detecting). This high degree of concurrency arises from the gateway’s role as a nexus for a multi-tenant environment, where countless independent application sessions are active simultaneously. (4) Traffic exhibits a highly skewed flow size distribution (pasham2022graph), characterized by the presence of long-lived, high-volume elephant flows alongside a large number of short-lived mice flows. (5) Processing logic is relatively simple and fixed (springCloudGateway), primarily driven by recurring functions such as ACL checks, metering, and next-hop determination, rather than arbitrary per-flow computations. It is also largely stateless in nature, with most flows requiring only basic forwarding or simple policy enforcement.
Design goals of our cloud gateway. Driven by these characteristics, we crystallize three fundamental design goals for our cloud gateway:
  • Low latency forwarding. To directly satisfy latency-sensitive traffic demands, the cloud gateway must minimize end-to-end transmission delays to ensure that these applications consistently meet user experience requirements. This goal is critically challenging in cloud networks due to deep topologies inherent to hyperscale deployments, where the co-existence of high throughput and million-flow tables compounds latency-critical path conflicts.
  • High throughput and large table support. To sustain such high throughput in ByteDance’s datacenters, the cloud gateway must deliver high-throughput packet forwarding and processing, while simultaneously maintaining massive routing and policy tables to support tens of thousands of tenants in a multi-tenant environment. Meeting both requirements simultaneously is challenging: petabit-per-second aggreate traffic places significant demands on gateway bandwidth, whereas large tables impose heavy pressure on memory resources.
  • Cost efficiency. At Petabit scale, even minor inefficiencies quickly translate into significant total cost. Cost efficiency necessitates co-optimizing procurement (CapEx), operational expenses (OpEx), and development velocity through judicious selection of forwarding components and architectures for minimal TCO. Because real-world traffic exhibits highly skewed flow size distributions and the required packet processing and forwarding logic is fundamentally simple, hardware caching and offloading represent a strategically sound design choice for cost efficiency. However, realizing hardware caching and offloading while maintaining flexibility and sufficient resources remains a non-trivial challenge.
2.2. Evolution towards Zephyrus
The Volcengine cloud gateway has experienced two major production-scale iterations. The evolution of Zephyrus is a story of systematically identifying and overcoming architectural bottlenecks. We first moved from a flexible but slow software gateway to a performant switching ASIC-based design(with x86 software assistance). However, new bottlenecks soon emerged, particularly the limited on-chip resources of switching ASICs. These constraints ultimately motivated our transition to a heterogeneous compositional architecture, which fundamentally overcomes the limitations of previous generation gateways and simultaneously achieves low latency, high throughput, rapid development, and cost efficiency.
The first generation: Software-only cloud gateways. Our initial cloud gateway adopted a fully software-based architecture, which is flexible and versatile but ultimately limited in performance and scalability (Kohler:Click:SOSP1999; Dobrescu:RouteBricks:SOSP2009; Han:PacketShader:SIGCOMM2010). This choice, driven by the need for rapid feature evolution under an MVP strategy, lets us leverage mature user-space packet I/O ecosystems (e.g., DPDK (dpdk_on_dpu), netmap (Rizzo:netmap:ATC2012), and IX (Belay:IX:OSDI2014)) for fast development. However, it embodied a fundamental contradiction between general-purpose CPUs for diverse computations and the repetitive, structurally simple work of packet forwarding.
As traffic grew, tenant performance isolation became critical. While a default per-attachment cap (e.g., 5 Gbps) effectively contained noisy neighbors, this rigid approach inherently conflicted with east-west traffic demands, resulting in non-negligible packet drops and application degradation. To mitigate this, we explored edge-based congestion-aware limiters (e.g., EyeQ (EyeQ-NSDI13)). However, their susceptibility to non-stationary signals and delayed feedback induces oscillatory throttling under bursts, degrading tail latency by a factor of 1.7–4.9 in production (Protego-NSDI23). Perhaps the most critical bottleneck, even with horizontal scaling, is the elephant flow problem. Hash-based load balancers, from in-network ECMP to host-level RSS, deterministically map large flows to a single CPU core. This creates severe hotspots that fundamentally limit the entire system’s throughput to that of its single most overloaded core (Hedera-NSDI10). This per-core bottleneck is a fundamental barrier to scalability, as simply adding more servers fails to address the underlying traffic skew.
These challenges are evidence that general-purpose CPUs are ill-suited for the simple, repetitive work of line-rate packet processing. Although initially flexible, a software-only gateway presents a fundamental barrier to achieving scalable, high-performance operation.
The second generation: Tofino-based gateway with x86 CPU assist. To address the challenges inherent to software-based gateways, we designed Tofino-based cloud gateways with the assist of x86 CPUs. Using programmable ASICs, our cloud gateways achieve line-rate performance, delivering 1.6 Tbps of throughput per device while maintaining a consistent ultra-low latency of 2µs. However, even after integrating four independent pipelines into a unified processing unit, the Tofino ASIC suffers from critically scarce on-chip resources. These constraints make supporting large-scale table entries and somewhat complex processing logic extremely challenging. As our tenants continue to grow, the on-chip SRAM required to store massive routing tables (most notably the VM-to-Node-Controller mappings) becomes nearly exhausted, despite applying optimizations including compiler-driven resource packing and key-compression for IPv6 entries(pan2021sailfish). Table 2 details the utilization of SRAM in twelve consecutive pipeline stages within the fourth pipeline where the P4 compiler maps the VM-NC tables, a position rigidly dictated by the data flow dependencies of the previous processing stages.
Table 2. SRAM Utilization Hotspot in Pipeline3. The extremely high average of 97.4% across 12 stages, with five individual stages hitting the 100.0% physical limit, reveals a severe and unavoidable localized resource bottleneck.
Metric S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11 S12 Average
SRAM Util. 100.0% 100.0% 100.0% 97.5% 98.8% 98.8% 98.8% 100.0% 98.8% 100.0% 98.8% 77.5% 97.4%
In addition to localized on-chip memory shortage, offloading the data plane to Tofinos significantly reduces development flexibility. Some complex operations are difficult to implement efficiently on Tofino pipelines (e.g., NAT, sophisticated Layer-4 load balancing) . Furthermore, computationally intensive functions such as IPsec encryption or operations requiring large elastic buffers, such as IP fragment reassembly, are fundamentally ill-suited for the relatively simple and fixed processing model of Tofinos, forcing these tasks back to software.
Towards the third generation: hybrid programmable gateways based on heterogeneous hardware. The limitations of previous generations motivate the need for a new class of hardware that bridges the gap between CPUs and Tofino ASICs. Rather than replacing existing components, this hardware is designed to work in concert with Tofino and software, forming an integrated next-generation gateway. Such hardware must:
  • Provide abundant on-chip resources for table entries.
  • Maintain programmability and flexibility.
  • Achieve competitive throughput and latency.
We conducted extensive analysis and experimentation to evaluate two potential options that could meet these requirements: FPGAs and DPUs. We ultimately selected DPUs, as they offer several advantages over FPGAs in practice:
  • Lower power consumption and cost. Cloud gateways with DPUs consume about 21% less power and reduce overall procurement cost by 14% compared to FPGA-integrated designs. When scaled to large deployments, these savings translate directly into lower CapEx and OpEx through reduced hardware spend and energy usage.
  • Lower development complexity and faster deployment. Compared to FPGAs, DPUs are significantly easier to develop with. DPU programming typically relies on high-level languages such as C/C++ and P4, supported by mature software ecosystems and toolchains (dpdk_on_dpu; nvidia_dpu_whitepaper). In contrast, FPGA development requires low-level hardware description languages (e.g.,Verilog / VHDL) and involves long synthesis and verification cycles (xilinx_hdl). These differences make DPUs far more developer-friendly.
  • Easier software access. DPUs integrate on-board ARM cores, enabling native DPDK deployment without external dependencies, as the ARM execution environment supports direct DPDK execution (nvidia_dpu_whitepaper; dpdk_on_dpu). In contrast, most FPGAs rely on host CPUs via PCIe interconnects, which require custom drivers and introduce complexity for DPDK integration (xilinx_smartnic). This architectural difference simplifies DPDK implementation on DPUs by eliminating host-dependent communication layers.
While DPUs provide compelling benefits, deploying them in coordination with switching ASICs in cloud gateways is a non-trivial task. With no prior blueprint to follow, our design of Zephyrus required rethinking the gateway architecture from the ground up. We systematically addressed these challenges by establishing guiding principles.
3. Zephyrus’s Architecture
This section first establishes the guiding principles for our design and then presents the architecture of Zephyrus. Detailed processing logic and optimization techniques are deferred to Section 4.
3.1. Guiding Principles
To navigate the complexity of the system, we established a set of guiding principles before our design.
Execute each processing stage on the most suitable hardware. The division of labor between switching ASICs and DPUs is informed by their intrinsic hardware characteristics. Switching ASICs are tasked with high-throughput, low-complexity processing; they can accommodate millions of simple table entries and process them with ultralow latency. Moreover, the relatively ample TCAM resources on switching ASICs motivate us to perform LPM matching on the ASICs, while delegating exact matches to DPUs.
Hierarchical offloading architecture. Hierarchical offloading is crucial in heterogeneous systems, as components exhibit an inherent programmability–performance gradient: switching ASICs offer the least programmability but the highest performance, DPU pipelines sit in the middle, and software offers the most programmability but the lowest performance. This naturally extends the conventional two-tier fast/ slow path model to a three-tier processing hierarchy: the ASIC pipeline (Tier 1), the embedded pipeline of the DPU (Tier 2) and the general-purpose ARM cores of the DPU (Tier 3). This structure recursively applies the fast/slow path paradigm: Switching ASIC and DPU pipelines form layered, high-speed mechanisms, while ARM cores provide the ultimate policy decisions.
3.2. High-Level Architecture
Figure 3. Three typical datapaths for different types of traffic
The end-to-end packet-processing pipeline is partitioned into three coarse-grained stages: pre-DPU (Tofino), DPU, and post-DPU (Tofino). Upon pipeline entry, each packet is tagged with a Service Identifier (SvcID) for packet classification and subsequent processing. All packets traverse Tofino pipelines (pre-DPU and post-DPU), but only a subset of packets require DPU processing. Within the DPU, the packets are further divided into Software Slow Path or Hybrid Fast Path for accelerated processing. Figure 3 illustrates three typical datapaths for different types of traffic in our hierarchically co-offloading architecture.
ASIC-Only Path (the red line). For traffic not destined for VMs within the gateway’s local datacenter (e.g., VM-to-Cross-Region-VM, VM-to-IDC, VM-to-Internet), the DPU stage is bypassed. Packets on this path are marked accordingly at the end of the pre-DPU pipeline to indicate that the packet will bypass the DPU stage, processed entirely within the switching ASIC. This path achieves the lowest latency for minimal processing, leveraging relatively fixed logic in the switching ASIC.
Hybrid Fast Path (the blue line). For traffic destined to local VMs, due to limited memory in switching ASICs, it is necessary to add additional processing logic and storing entries in DPU SRAM. Packets on this path propagate metadata from the pre-DPU stage to the DPU stage for exact-match lookups, where a flow cache (stored in DPU DRAM) performs accelerated matching. Cache-hit packets are processed by the post-DPU pipeline, whereas missed packets are handled via the Software Slow Path on ARM cores.
Software Slow Path (the purple line). Upon a cache miss for local VM-destined traffic, packets are punted to the DPU’s ARM cores for software processing. The flow offloading agent subsequently installs a forwarding rule in the hardware cache, offloading future packets of the same 5-tuple. Provided that the rule persists in the cache(without eviction due to memory constraints or flow reconfiguration), subsequent packets with matching 5-tuples will be offloaded to Hybrid Fast Path.
Figure 4. Detailed pipeline organization.
Figure 5. Table lookup sequence in Tofino pipelines.
4. Compositional P4 Pipeline Design
Beyond fundamental cloud gateway functions (e.g., ACL, NAT, AZ affinity, mirroring, and metering), we developed targeted optimizations to improve performance. In this section, we first describe the detailed architecture of the unified P4 pipeline, followed by detailed packet processing, then present key mechanisms enabling efficient, consistent, and scalable packet processing in cloud gateways.
4.1. Detailed Pipeline Organization
Figure 4 shows the pipeline organization in detail; protocol and control packets are diverted at ingress pipe 3, as indicated by the red dashed arrows. Given the limited resources of a single Tofino pipeline, we evaluated the required memory footprint and fine-grained pipeline-stage count for the major tables (e.g., interface tables, metering tables, routing and policy-routing tables, and next-hop determination tables) and packet-processing logic, concluding that four independent Tofino pipelines are needed to complete the tasks. To this end, we apply pipeline folding to assemble the four independent pipelines into a single folded pipeline that traffic traverses. After folding, we quadruple the available Tofino ASIC resources at the cost of reducing throughput from 6.4 to 1.6 Tbps (4× reduction). We then employ four 400 Gbps embedded DPUs in parallel to match the 1.6 Tbps line rate of the folded pipeline, using ECMP to distribute traffic across the DPUs.
We also evaluated an alternative design where the DPUs are placed serially after the entire folded Tofino pipeline. Although this arrangement might appear to simplify management, it introduces two critical drawbacks. First and most fundamentally, it misaligns the roles of the hardware. This design forces DPUs to handle line-rate forwarding, a task for which they are not optimized. DPUs excel at offloading complex operational tasks (e.g., encryption, deep packet inspection), not replacing the raw forwarding power of a switching ASIC (ermerins2021bluefield). Second, the alternative design suffers from a severe bandwidth penalty. This forces a hard partitioning of the front-panel interfaces that half of the ports to be consumed by the internal ASIC-DPU connection and the other half for external connectivity. Consequently, the bandwidth is halved from 1.6 Tbps to 800 Gbps. The key advantage of the original design is that traffic enters and exits the DPU from the same Tofino pipeline, achieving full-duplex communication without bandwidth loss, and it also places the DPU exactly where complex operations are needed. By comparison, LuoShen sacrifices a quarter of its external ports to connect with the CPU and FPGA (pan2024luoshen), while the current design avoids any bandwidth loss by allowing data to enter and exit the DPU from the same Tofino pipeline. Therefore, we adopt the three-stage ”sandwich” architecture for our design.
Figure 6. DPU packet processing pipeline with fast/slow path interaction
4.2. Packet Processing across Heterogeneous Hardware
Guided by our Principles discussed in Section 3, our design strategically maps different table types to the hardware components that are best suited for them. Given the scarcity of TCAM on DPUs but their abundant SRAM and DRAM, we place large-scale exact-match tables (e.g., for VM-destined traffic) on them. In contrast, Tofinos are equipped with TCAM, which is particularly well-suited for longest-prefix match (LPM). For protocol and control packets, DPU involvement is unnecessary. Allowing them to traverse the whole pipeline not only lengthens the processing path but also introduces potential risks to stability. To address this, we divert such traffic at the ingress pipe 3, bypassing the DPU and forwarding directly to the CPU. Prior to CPU handoff, essential protocol-specific handling (e.g., BGP processing) is performed at the ingress stage, ensuring correctness since certain BGP functions depend on table lookups. This selective diversion avoids unnecessary traversal of heterogeneous hardware, reduces latency, and guarantees that control-plane traffic follows the shortest and most stable path.
Table lookups in Tofino pipelines. As illustrated in Figure 5, Tofino executes packet processing as a sequence of table lookups. Incoming packets first look up a traffic classification table, which recognizes two categories of packets: VXLAN-encapsulated tenant traffic (identified by remote VTEP’s physical IP) and protocol/control packets. Tenant traffic passes through a tenant identification table that determines the policy routing table to look up. At this point, the system translates the VXLAN Network Identifier (VNI), a generic overlay identifier, into a policy routing identifier, which selects the appropriate virtual routing context. This translation is key to achieving resource isolation, since each tenant is associated with a dedicated policy routing table.
After the tenant context is established, packets proceed to the policy routing table, which supports multiple routing instances as commonly deployed in public cloud environments. Within a single tenant (i.e., the same VNI), multiple subnets or security domains may coexist. Consequently, the inner virtual source IP address is used to select the appropriate routing instance. Packets then enter the routing table, which performs a lookup on the inner destination IP address to obtain either a nexthop index or an ECMP group index. ECMP selects a nexthop based on the five-tuple hash, providing load balancing across multiple equal-cost paths. Finally, the nexthop table provides the precise forwarding information, including the output port and encapsulation.
Packet processing in DPUs. Once a packet enters the DPU, it first traverses our P4-programmable fast path. The programmable pipelines, p4Ingress and p4Egress, implement flow and action logic across eight-stage match–action (MA) tables. These stages cover packet parsing, classification, flow lookup, and forwarding decisions. Fixed-function modules, including P4RxDMA and P4TxDMA, handle packet DMA operations and integrate vendor-customized lookup.
As illustrated in Figure 6, for cache miss or failed version check (expired packets, detailed in 4.3), we redirect them to the slow path, where the ARM cores running the DPDK-based gateway program reexecute policy enforcement, ACL checks, and nexthop computation. The results are then synchronized back into the fast path, ensuring that subsequent packets of the same flow can be processed entirely in hardware.
Table lookups in DPUs. When a packet enters the DPU pipeline, it first performs a lookup in the Flow Table, which primarily returns the nexthop index (pointing to the precise nexthop entry in Tofino at the post-DPU stage) and the version number (detailed in 4.3). The VM–NC (VM-to-Node Control) table maintains the mapping between virtual machines and their hosting nodes, allowing the system to efficiently determine the physical host associated with a given VM. Since the VM–NC table is the largest lookup structure and previously caused localized hotspots in Tofino, it has been relocated to DPU DRAM. On the software slow path, the system leverages the routing table identifier embedded in the packet metadata (obtained during the Tofino pre-DPU stage) as an index to selectively retrieve nexthop and VM–NC entries as needed.
4.3. Mechanisms for Consistent and Efficient Heterogeneous P4 Pipeline
Hash-table coalescing on DPUs. On DPUs, each additional hash table introduces extra processing logic, including key extraction, hash computation, memory lookup, and action execution. DPUs provide only limited TCAM capacity, which is insufficient for efficient longest-prefix matching (LPM). As a result, they primarily depend on SRAM/DRAM with hash indexing, favoring exact-match lookups but performing poorly on LPM. As the number of tables grows, the pipeline accumulates more control branches, forcing the DPU to schedule a larger set of action executions and then amplifying processing overhead, so that we need to try our best to coalesce our hash table. We observe that due to the specific traffic characteristics of cloud gateways, instances of multiple matches in longest-prefix matching (LPM) are rare. When transforming LPM into exact-match lookups, each candidate prefix must be probed in sequence, and each probe requires a base hash table and an overflow hash structure to handle collisions. In order to mitigate overhead, we perform collision checks in the control plane prior to rule installation. If no conflicts are detected, the entries are consolidated into a single table. Furthermore, we can guide user configurations to better align with our hash-coalescing scheme. In the absence of multiple matches, hash-table coalescing substantially reduces both processing complexity and resource consumption.
Minimal yet expressive metadata for communication. Since our DPU provides only eight P4 pipeline stages, we delegate the identification lookups to Tofino pipelines and carry the result to the DPU through metadata. We carefully design the metadata structure (illustrated in Figure 6) so that it conveys only essential information. To simplify processing, we align the header length for packets sent from egress pipe 2 to DPUs and for packets returned from DPUs to ingress pipe 2, avoiding the need for DPDK to repeatedly insert or remove headers. For the private protocol design, our initial format was Ethernet + DPU key + user Ethernet, but this proved cumbersome for each heterogeneous component to process. We later switched to Ethernet + private key + user payload, reusing the original Ethernet header and restoring it only when the packet leaves the system.
Version-based flow update mechanism. Flow updates are particularly challenging in heterogeneous hardware environments. Forwarding rules are distributed across Tofino ASICs, DPUs, and CPUs, each with distinct data structures and update mechanisms. Synchronizing these views consistently is non-trivial, especially given the mismatch in update timescales. Our solution associates each instance with a unique service ID, assigned in the Tofino ASIC and carried in the metadata to the DPU pipeline. Each service object is also tagged with a version number that reflects the latest configuration. Upon reconfiguration, the controller in control plane updates the version number. As illustrated in Figure 6, the DPU pipeline validates each flow by comparing its version with the current service version; a mismatch redirects the packet to the slow path for recomputation. This design provides a lightweight yet effective mechanism for maintaining consistency during flow updates, requiring only a version check in the fast path.
5. Evaluation
In this section, we evaluate the performance of Zephyrus in a realistic experimental setting. We show that:
  • Compared with the software-only cloud gateway, Zephyrus delivers lower latency and zero packet loss (Section 5.1).
  • Compared with the Tofino-based cloud gateway, Zephyrus offers more flexible on-chip resource utilization and supports substantially larger table capacities (Section 5.2).
5.1. Performance Comparison
(a) Packet Loss Rate
(b) Average Latency
(c) P99 Latency
Figure 7. Performance comparison of three cloud gateway versions
Our testbed comprises a Top-of-Rack (ToR) switch, two virtual machines (VMs), and multiple versions of a cloud gateway. The ToR provides intra-rack connectivity, the VMs act as traffic source and sink, and the cloud gateway sits in the forwarding path between them. On VM1, we run the DPDK testpmd application to generate traffic at a fixed rate of 10 Mpps. On VM2, we run a DPDK-based forwarder that performs a five-tuple swap (i.e., it swaps source/destination IP addresses and ports, and updates L2 headers as needed) and transmits the packet back toward VM1. As a result, each packet traverses VM1 $\rightarrow$ ToR $\rightarrow$ gateway $\rightarrow$ ToR $\rightarrow$ VM2 $\rightarrow$ ToR $\rightarrow$ gateway $\rightarrow$ ToR $\rightarrow$ VM1, forming a round trip, and the cloud gateway processes packets with 20Mpps. We measure packet loss and end-to-end latency (round-trip time, RTT) to evaluate the different gateway designs.
Packet loss. We sweep the total packet count per trial from 1K to 10M (at a fixed sending rate of 10 Mpps) and record the number of packets lost at each setting (Figure 7(a)). Both the Tofino-based gateway and Zephyrus exhibit zero packet loss across the entire sweep; when sending 10M packets, not a single packet is lost. In contrast, the software-only gateway starts dropping packets once the total packet count exceeds 5M; at 10M packets, the loss rate exceeds 5%. Such packet loss rate degrades the gateway’s delivered performance, harms service quality, thus is unacceptable in real-world deployments.
(a) 1K
(b) 1M
(c) 4M
Figure 8. Latency distribution with different packet numbers
(a) Pipeline 0
(b) Pipeline 1
(c) Pipeline 2
(d) Pipeline 3
Figure 9. PHV usage across pipelines (the worst case)
(a) Pipeline 0
(b) Pipeline 1
(c) Pipeline 2
(d) Pipeline 3
Figure 10. Total on-chip resource usage across pipelines
Latency. We evaluate latency within the no-loss regime (1K–4M packets). The results (Figures 7(b)7(c)) show that Zephyrus substantially reduces both average and P99 latency compared to the software-only gateway: at 4M packets, the average RTT decreases by 68.7% and the P99 RTT by 79.3%. Compared with the Tofino-based gateway, Zephyrus incurs only a modest additional delay of about 8µs, and both designs meet the requirements of cloud–edge services. The latency distribution of three gateway versions in Figure 8 further show that the Tofino-based gateway and Zephyrus maintain relatively stable latency, whereas the software-only gateway exhibits pronounced variability and steadily increasing latency with load. The long-tail latency distribution of software-only gateway has a detrimental influence on QoS and user experience (e.g., timeouts and jitter-sensitive degradation), making performance improvements necessary for production deployments.
(a) Pipeline 0
(b) Pipeline 1
(c) Pipeline 2
(d) Pipeline 3
Figure 11. SRAM usage over stages across pipelines
Figure 12. SRAM saturation percentage across pipelines
Figure 13. Table capacity of different versions
5.2. Resource Usage
PHV utilization. Figure 9 compares PHV utilization for the Tofino-based cloud gateway and Zephyrus. The results show that the 32-bit PHV constrains the scalability of the Tofino-based gateway: in the worst case, pipeline 0 nearly exhausts its 32-bit PHV capacity. In contrast, Zephyrus preserves ample headroom across all three PHV widths (8/16/32-bit) and pipelines; in every pipeline, utilization for each width remains below 85%, enabling further scaling.
On-chip memory utilization. Figure 10 summarizes on-chip resource utilization for the Tofino-based gateway and Zephyrus. Among these resources, SRAM is the primary bottleneck. As shown in Figure 11, the SRAM in pipeline 3 is nearly exhausted on the Tofino-based gateway, constraining its further scaling; in contrast, Zephyrus distributes SRAM more evenly across all four pipelines. Figure 13 further quantifies saturation as the per-pipeline percentage of stages with $95\%$ SRAM utilization. For the Tofino-based gateway, only 1 out of 12 stages in pipeline 3 remains unsaturated, whereas for Zephyrus the fraction of saturated stages stays below 60% in every pipeline, providing a substantial margin for growth and reducing the risk of resource scarcity.
[…]
In general, Zephyrus delivers comparable little-loss performance to the Tofino-based gateway while substantially improving scalability and functionality, and it markedly outperforms the software-only gateway in both packet loss and latency.
6. Related Work
In recent years, Data Processing Units (DPUs) have rapidly emerged as a key enabler for cloud and datacenter infrastructure. Recent work highlights their diverse capabilities, including storage offloading (zhao2024tianmen), near-storage secure filtering (batsoyol2025dpuf), and 6G user-plane acceleration (lee2024upf), demonstrating their growing role in different fields. Sirius (bansal2023disaggregating) further demonstrates this trend by disaggregating a broad class of stateful network functions into pools of P4-programmable NIC cards.
In the context of cloud gateways, previous state-of-the-art works primarily focus on using FPGAs (pan2024luoshen; lu2025albatross) and/or switching ASICs (pan2021sailfish; pan2024luoshen) for acceleration. Sailfish (pan2021sailfish) uses Tofino ASICs to accelerate forwarding, employing a series of hardware and software co-designs to address the on-chip memory limitations of the Tofino chip. LuoShen (pan2024luoshen) leverages FPGAs as an intermediate layer between CPUs and Tofinos to handle complex high-bandwidth forwarding, achieving 1.2 Tbps throughput. Albatross (lu2025albatross), in contrast, is based on FPGAs and x86 CPUs, delivering FPGA-based packet-level load balancing to the host without a switching ASIC.
In contrast, as a centralized gateway, Zephyrus adopts DPUs to augment the Tofino pipeline. This design significantly expands effective table capacity, overcoming the resource shortage limitations in architectures such as Sailfish and Luoshen. Meanwhile, Zephyrus’s three-stage ”sandwich” architecture improves its throughput to 1.6 Tbps, a 33% improvement over Luoshen and a 100% improvement over Albatross. To overcome memory limitations, Albatross, which is based on FPGA, suffers a significant performance regression in throughput (4x), packet rate (18x) and latency (10x) compared to Tofino-based gateways. By comparison, Zephyrus uses a DPU-augmented architecture to overcome resource limitations; by hierarchically co-offloading over 99% of packets to hardware, it achieves only a slight performance loss compared to Tofino-based gateways.
As cloud workloads diversify and traffic scales non-linearly, a growing number of hyperscalers are exploring heterogeneity in the network dataplane (Catapult2018). However, this very heterogeneity introduces significant development and operational complexity. Programming these disparate hardware targets requires specialized teams and fragmented toolchains, leading to long and costly development cycles. We introduce a unified P4 pipeline to accelerate development cycles (Bosshart:P4:2014), significantly simplifying the development and operational complexity.
7. Experiences
7.1. Development and Deployment Experiences
Heterogeneous Hardware Development. Managing a heterogeneous data plane presents significant operational challenges. To address these, first, we implemented a flexible packet distribution strategy. By default, hash-based routing balances traffic between the Tofino switch and the DPU processing nodes. However, recognizing that hashing can suffer from traffic skew, we provide a dynamic one-to-one mapping mode to deterministically route specific flows, mitigating hotspots and preventing node overload. Second, to provide fine-grained visibility into this complex pipeline, we implemented a packet-level tracing mechanism. Each packet is marked with a unique identifier at ingress, enabling the precise tracing as it traverses the Tofino, the DPU, and other processing stages. This replaces what is typically a complex and time-consuming debugging task with a process of systematic analysis. As a result, operators can quickly identify the exact stage of failure, streamlining fault diagnosis in a heterogeneous environment.
System Upgrade. We upgraded the system through multiple major releases using a staged deployment approach. To minimize downtime during transitions, we routed traffic to x86‑based fallback paths, enabling rapid rollback when necessary. The rollouts were carried out in phases: internal pilot groups, internal key operational teams, external users, and finally all users of the legacy system. External rollouts followed the same progression (first pilot deployments, then expansion to key external partners). Central to this strategy was continuous, real-time monitoring. This enabled our teams to detect and resolve any performance regressions or anomalies prior to each subsequent deployment phase, ensuring our service-level objectives (SLOs) for latency and availability were consistently met.
Business Expansion. In some deployment scenarios, a single data center must carry aggregate bandwidth at the TB level. During business expansion, additional capacity is typically provisioned before traffic is routed onto new resources. However, when traffic is migrated in batches after expansion, newly added nodes may not align with the existing traffic-distribution layer or routing symmetry; consequently, individual links can transiently receive more than half of the load during the transition, risking node overload or failure. To mitigate this, we conduct a capacity and topology assessment prior to expansions. If the assessment reveals a high risk of transient traffic spikes, we adopt a more cautious, two-step transition plan. First, traffic is temporarily redirected to underlying physical switches to ensure a more uniform load distribution. Then, once fully prepared, all new nodes are brought online simultaneously in a single, coordinated update.
Disaster Recovery. Our disaster recovery strategy is designed for rapid fault detection and recovery across multiple failure domains. When a link fails, the underlying routing protocol automatically detects the status change and isolates the link, preventing fault propagation. For a device-level failure, traffic is immediately rerouted to a standby device based on pre-established connectivity, preserving service continuity. Primary and backup clusters are deployed in separate data centers with segmented routing to flexibly schedule traffic, allowing rapid switching to backup sites in case of datacenter-level failures and thereby ensuring high availability. These measures enable swift, appropriate responses to faults, minimizing business impact and maintaining system stability.
7.2. Lessons Learned
Upgrade the system selective. During the upgrade process, we selectively upgraded only the necessary components. The previous Tofino-based cloud gateways remain well-suited for fixed-function, lightweight forwarding tasks, so we retained parts of this infrastructure (e.g., IGW, CEN) in Tofino-based gateways. This separation reduces Zephyrus’s resource footprint, lowers the upgrade burden, and, in turn, significantly reduces costs while enabling rapid deployment.
Choose the right programming language. We find P4 to be the ideal language to describe gateway forwarding logic, as its declarative style naturally expresses the packet processing pipeline (parser, match action, deparser) and compiles directly to fixed ASIC stages with stable latency. Compared to Verilog/VHDL, which imposes a steep learning curve and long development cycles, P4 achieves much higher development efficiency and benefits from a mature ecosystem (e.g., bmv2, P4Runtime), making it a natural fit for cloud gateway development.
Conduct thorough evaluation before development. We first evaluated our growth in production traffic and recognized the need to redesign the next generation gateway. Through a systematic evaluation of hardware options, we selected DPUs over FPGAs. While both provide hardware acceleration, DPUs offer significantly lower development complexity. Prior work shows that FPGA-based data plane designs often require tens of thousands of lines of Verilog or HLS code for packet parsing and control logic (e.g., FpgaNIC (fpganic_atc22), NetFPGA (netfpga_p4)), while DPU development leverages C/C++ with DPDK libraries or P4 for specific modules, typically implemented in only a few thousand lines of code (dpdk_doca; dpdk_on_dpu). Guided by workload analysis including traffic composition, frequency distributions, and resource demands, we adopted a hierarchical forwarding architecture. We also evaluated the expertise within our team and assigned responsibilities accordingly before starting implementation. This careful upfront evaluation directly contributed to reduced TCO, and substantially fewer man-months compared to a less structured development process.
Make the system future-proof. Fueled by continuous breakthroughs in SerDes (Serializer/Deserializer) technology, the aggregate throughput of switching ASICs has followed an exponential growth trajectory, doubling approximately every two years. We implemented and deployed Zephyrus using Intel Tofino switches, which were the state-of-the-art P4-programmable ASICs when our project began. As a long-running system, we must anticipate a future where our existing inventory of a specific ASIC model becomes depleted during large-scale deployment, and the broader hardware landscape will have evolved. Designing an abstract architecture decoupled from specific hardware, building on hardware types that constantly advance, and maintaining an experienced developing team enable us to rapidly migrate to more performant or cost-effective switching ASICs (e.g., Xsight Labs Unveils X1, Broadcom Tomahawk) with minor redevelopment. Meanwhile, in terms of DPUs, the Intel roadmap projects that its upcoming IPUs (a DPU equivalent design) will reach 800 Gbps in 2025 (kennedy2022ipu), offering strong evidence in favor of our design and showing that Zephyrus is well positioned for future evolution.
7.3. Broader Design Principles and Insights
The design, implementation, and deployment of Zephyrus have yielded insights that extend beyond our specific system. Here, we discuss the broader design principles and insights that have emerged from our work.
The principle of software-hardware asymmetry. Our experiences with Zephyrus suggest that architectures can benefit from an intentional software-hardware asymmetry. This experience indicates that hardware’s role is often best suited for low-latency, high-throughput, and repetitive data-plane tasks (e.g., packet parsing, encryption, and basic forwarding). In contrast, software appears well suited for managing complexity and evolving business logic (e.g., policy orchestration, cross-domain optimization, and anomaly detection (openflow08; b4_13; sonata18)). Zephyrus’s architecture is a direct embodiment of this principle, combining the Tbps-scale throughput of its ASIC data plane with the rich flexibility of its software components.
An explicit line between standardization and customization. Running a system like Zephyrus at scale forces us to confront a classic tension: how do we keep the system stable while constantly adding new features for different customers? Our answer was to draw an explicit line between standardization and customization. In our observation, core infrastructure components, such as routing protocols, hardware interfaces, and baseline control logic, benefit from conforming to industry standards to promote interoperability (Saltzer1984) and stability. Programmability, in turn, is leveraged at the service edge to enable fine-grained, tenant-specific features (e.g., differentiated QoS policies, dynamic filtering rules). This separation allows us to satisfy unique requirements while minimizing impact on the standardized core, thereby preserving system-wide consistency and robustness.
8. Conclusion
Zephyrus is ByteDance’s third-generation cloud gateway. DPU integration, together with the hierarchical co-offloading strategy, overcomes challenges and limitations of previous gateways, without compromising their core strengths. Built for the petabit era, Zephyrus scales to support massive traffic workloads while meeting emerging business needs. We are continuously optimizing this new architecture to deliver higher performance and meet emerging business demands. As we move forward with the system’s continuous upgrades and deployments, we are committed to sharing our operational experiences and insights with the cloud networking community.